# A Survey of Embodied AI in Healthcare: Techniques, Applications, and Opportunities

### v1

@@ Abstract
{IEEEkeywords}
embodied artificial intelligence; multi-modality; healthcare; large language model; world model
@@ 2.1.2 Interactive Perception
Figure 3: A detailed breakdown of EmAI core functionalities, highlighting the key approaches that support the functionality implementation.
2.2 Low-level Actuation
@@ 2.3 High-level Planning
Planning with end-to-end Embodied Large Models. These algorithms stand out by directly mapping high-level instructions to low-level actions, seamlessly integrating perception, planning, and control into a unified system. Recent research shows that these systems leverage deep reinforcement learning and imitation learning to streamline planning and decision-making in complex environments, often surpassing modular pipelines in adaptability and robustness [310, 311]. Notable frameworks such as SayCan [283], PaLM-E [284], and EmbodiedGPT [287] are designed to combine vision encoder embeddings with planning data from LLMs, directly informing the robotic policies for immediate action. However, training embodied end-to-end systems typically requires large-scale datasets. To address this need, simulation data is widely used for its efficiency and safety, though challenges arise from discrepancies in physics, sensors, and real-world complexity. Strategies such as domain randomization [312], domain adaptation [313], and hybrid approaches [314] have been developed to enhance adaptability and bridge the sim-to-real gap, improving real-world performance. In addition, it is also important to design benchmark to evaluate the embodied decision making capability among different LLMs [315].
@@ 3.1.1 Pre-Intervention Stage
Recent improvements in EmAI-related technologies for pre-intervention diagnostics and assessments are shaping a new AI-clinician collaboration in the intelligent hospital [356, 357, 358]. EmAI plays various roles in this context, reducing clinicians’ workload and accelerating diagnostic workflow.
Virtual triage nurse. In modern smart healthcare systems, EmAI-based virtual triage nurses replace human nurses play a pivotal role in streamlining patient management by directing individuals to the most appropriate clinical departments. These EmAI systems analyzed patient-reported symptoms and conduct department ranking based on symptom descriptions [359, 360, 361]. Beyond symptom-based sorting, advanced triage systems integrate wearable health data and EHR to provide a holistic assessment of conditions [362, 363, 364]. In emergency care settings, these systems can even predict patient outcomes and recommend intervention pathways, significantly shortening response times during critical situations [365, 366, 367, 368]. As healthcare systems become more interconnected, intelligent triage systems are increasingly serving as an efficient tool of future infrastructure [364, 369], enabling seamless coordination between primary care, specialist consultations, and hospital admissions.
@@
  • Radiology assistant. Interventional Radiology (IR) [419] focuses on using image-guided, minimally invasive techniques to perform diagnostic and therapeutic procedures. The integration of EmAI is further advancing IR by improving procedural accuracy, safety, and operator control in complex medical interventions. For exmaple, robotic systems have applied CT images to improve the accuracy of localization of lesion areas such as pulmonary nodules [107] and navigated placement of thoracolumbar pedicle screws [420] in both preoperative diagnosis and intraoperative procedures. Similarly, robotic systems leveraging X-ray imaging have enhanced cardiac interventions by enabling precise catheter navigation in endovascular catheterization [421, 422]. MRI-guided robotic interventions have facilitated resection margins targeting in procedures like tumor ablation, providing superior visualization of real-time tissue changes during operations to protect critical structures [423, 424, 425].
@@
In summary, fully autonomous ultrasound system [110], human-centric radiology assistant [433], and flexible robotic endoscope system [434] capable of navigating and adjusting diagnostic examination with minimal human supervision have been developed, achieving a higher level of intelligence and autonomy. Compared to human-led examinations, they can optimize workflows, improve patient care quality, and ensure a safer environment for all involved in the surgical process. Besides, these systems can also function as a module within a more comprehensive EmAI system such as surgical robots, showcasing opportunities to expand their applications.
@@ 3.1.2 In-Intervention Procedure
While the prior discussion centered around the planning of surgical skills into action-level details, within the contemporary medical landscape, there are complex multi-stage surgeries with long-term plans, dividing complex surgeries into multiple sessions for execution. Such surgeries often involve intricate coordination among surgical teams, precise management of patient health over extended periods, and adaptation to unforeseen intraoperative changes. Compared with traditional surgical planning methods, EmAI can enhance multi-stage planning for complex surgeries by integrating various system modules, such as diagnostic tools [395, 396, 458], real-time monitoring systems [459, 460, 461], and predictive analytics [22, 23] for long-term patient health management. These AI-driven modules facilitate dynamic adaptation of clinical intervention schedules and strategies based on assessment of patient conditions during successive surgeries [462, 463]. While significant advances have been achieved in individual EmAI modules, seamlessly integrating them into a fully functional surgical planner with global cognition remains a critical focus for future innovation.
@@
Surgical operator. Skillful operators are at the core of every surgical procedures, where precision and timely intervention are crucial. Challenges often arise due to the intricate nature of surgeries and the finite availability of expert operators, which can lead to bottlenecks in patient care. To enhance this, robotic surgery has been a promising trend to offer enhanced precision, dexterity, and minimally invasive approaches to common surgical procedures, leading to improved patient outcomes [464]. In healthcare, a dedicated robot is often more practical and reliable, and has been extensively studied in a variety of settings and shown to provide overall clinical benefits [465]. For example, in laparoscopic surgery, there are robotic systems autonomously performing intestinal anastomoses [466] and rectal cancer resection [467] under human oversight. In retinal microsurgery, comprehensive EmAI systems [468, 469] have been developed to integrate real-time surgical object tracking, segmentation, and model predictive control for intraoperative navigation in delicate microsurgical environments wherein precision is decisive.
@@
For general surgical skills, the potential of robotics has also been widely demonstrated. Robots driven by deep learning and active sensing strategies are trained to use needles through HOUSTON algorithm [470], which can localize, grasp, and hand over unmodified surgical needles to complete precise manipulation. Their good suturing skill that requires stable and delicate control system has also been proved by the SmartArm robotic system for a neonatal chest surgery [435]. Similarly, for thread detection and interactive perception, real-time robotic approach [437] is presented for autonomously shortening the tail of a surgical suture using a self-supervised learning framework. In complex surgical scenes, multiple EmAI subsystems may collaborate to integrate multi-source perception and support robotic surgery, requiring a robust control system to direct and coordinate their respective perception-cognition-communication-action loops [471]. Beyond visual perception, other modalities including kinematic data [472], audio data [473], language instructions [102], and tactile perception [474] are integrated to form a holistic view of real-world surgical environments.
@@
Patient digital twin. A patient digital twin represents a detailed, dynamic model of a patient’s biological systems or a part thereof (e.g., anatomy), created using comprehensive and accurate medical data such as imaging studies, physiological measurements, and diagnostic results [114]. It is often combined with advanced visualization devices (e.g., VR/AR) to provide interactive and observations, and surgeons can plan, simulate, and optimize surgical pathways based on them [508, 509]. AI-based digital twin creation technology facilitates the identification of surgical patterns [438], prediction of complications or procedural outcomes [509], and generation of medical reports [510, 511]. This technology is often employed to accelerate the learning curve of novice surgeons, helping them familiarize themselves with anatomical structures, surgical procedural contexts, and disease progression [512, 513, 514, 515, 516]. Furthermore, multiple downstream applications have been benefited from the surgery digitization process [517]. Clinicians can practice surgical skills and explore human anatomy without relying on real anatomical models using digitalized surgical platforms [518], overcoming the limitation of high-cost and scarce training samples. It also offers digital and accurate platform for robots to train on, improving safety and reliability before deployment in actual surgical procedures [519]. Another approach involves photorealistic surgical images [520] and videos [521] synthesis to benefit the training process of EmAI systems, mitigating challenges related to the high costs and ethical concerns of obtaining surgical data.
@@ 3.2.1 Assistive Robot
The integration of EmAI into healthcare has led to the development of assistive robots that enhance daily care and support of patients. This section focuses on three key applications including social assistance, daily living assistance, and mobility assistance.
Social guide. In contemporary society, individuals with Autism Spectrum Disorder (ASD), Bipolar Disorder, or other social challenges often face significant barriers in accessing adequate social support. The advancement of EmAI systems, especially those driven by LLMs, has facilitated the development of innovative solutions aimed at addressing these needs [583, 584]. The NAO robot and QTrobot have been utilized to enhance facial recognition and improve eye contact, serving as therapeutic mediators for young people with autism and helping them learn to interact with others [585, 586, 587, 588]. Through structured interactions, children can practice social skills such as imitation, turn-taking, engagement, and empathy [589]. Studies have indicated that children with ASD often respond positively to robot-assisted therapy, showing increased engagement and reduced anxiety during sessions [590, 591, 592]. Beyond psychological disorders, EmAI systems like ZORA robots [593, 594, 595] have assisted children with severe physical disabilities in achieving therapeutic and educational goals, as well as helping patients with dementia optimize communication. We present a pipeline for an EmAI system that supports children with social disorders, as shown in Figure 7.
@@ 3.3.1 Rescue Robot
EmAI powered robots are increasingly developed to serve as human assistants or replacements in various high-risk environments. In rescue scenarios, their potential has been widely explored due to their ability to operate safely in dangerous environments and urgent conditions. In the aftermath of natural disasters or life-threatening emergencies, search and rescue robots aim to provide rapid aid and emergency health services to victims. Thanks to the advancement of recent AI approaches including MLLMs and VLMs, EmAI-equipped robots can efficiently locate and assist survivors while supporting rescue operations [642, 643, 644, 645] in disaster-prone areas, such as earthquake zones. A representative workflow of EmAI-based rescue robots is illustrated in the Figure 8.
@@
The pandemic has also significantly increased the demand for service robots to replace human labor in contaminated areas, particularly for disinfection tasks. This has made efficient and rapid disinfection in hospitals crucial, sparking researches into disinfection robots. By integrating mobile robot platforms with advanced disinfection technologies, such as hydrogen peroxide atomization [675] and UVC devices [676, 677], and equipping them with EmAI navigation systems, these robots are able to automate cleaning in complex indoor environments. They are designed to disinfect both floor surfaces and air by eliminating harmful pathogens using a combination of vacuum cleaners, sanitizers, and UV light disinfection methods [678, 679, 680, 681]. By upholding rigorous hygiene standards, these robots significantly mitigate the risk of cross-contamination in healthcare settings.
@@ 3.4 Biomedical Research
EmAI also shows great potential in biomedical and healthcare research [682]. It can synergistically combine human creativity and expertise with AI’s capabilities in handling vast biomedical datasets, navigating complex hypothesis spaces, and executing repetitive experiments. The core functionality of EmAI systems in biomedical research involves planning discovery workflows, performing self-assessments to recognize and address knowledge gaps, and employing structured memory for continual learning [27, 80].
Automated lab technician. Experiments are critical parts in biomedical research. EmAI-driven robots automate repetitive tasks, increasing throughput and allowing researchers to focus on more complex analysis tasks. Scientists have created robot chemists capable of autonomously conducting complex chemical reactions and analyses [683, 684]. These robotic systems apply EmAI in laboratory automation, enabling them to design experiments [685, 686, 687], execute procedures [688, 689], and interpret results [690, 691, 692] without human intervention. For instance, a platform leverages AI-based synthesis planning, trained on millions of reactions, to propose synthetic routes, which are then fine-tuned by expert chemists and executed on a robotic flow system [693]. This setup allows for scalable, reproducible synthesis of a variety of compounds, such as pharmaceuticals. Another AI-driven robotic chemist also demonstrated its ability to autonomously perform tasks ranging from synthetic planning to executing experiments in batch reactors [694]. In the field of materials science for photo-catalytic hydrogen production, the robot operates autonomously in a laboratory environment, using Bayesian optimization to navigate a complex experimental space with ten variables, performing 688 experiments over eight days [683]. This workflow is illustrated in Figure 9, which shows a regular automation process of chemical experiments. In summary, EmAI can combine precision handling with autonomous decision-making, allowing it to run experiments continuously without human intervention. These experiments setups minimize human bias, accelerate discovery and production, and enable exploration of large experimental spaces.
@@
Pioneer in drug discovery. The traditional drug discovery process is typically time-consuming and involves substantial costs, presenting significant challenges for pharmaceutical research and development. However, the integration of EmAI into the drug discovery process not only shortens the overall development timeline but also reduces the associated expenses. AI-driven robotic systems can perform rapid screening of vast chemical libraries, identifying potential drug candidates more efficiently than manual methods [695]. In high-throughput screening (HTS) systems, physical components are primarily responsible for automating processes such as sample handling, reaction setup, and data collection through robotics and specialized software for instrument control [696]. By employing liquid handling devices, robotic technology, detectors, and specialized software, researchers have also achieved automation and high efficiency in HTS processes [697] to swiftly evaluate large libraries of compounds for activity against specific biological targets. Additionally, in the analysis and prediction of biological sequences, such as DNA, RNA, or protein structures, LLMs and other approaches are applied to decipher patterns in biological sequences [698]. By treating nucleotides and amino acids as words, they can predict biological structure and functions, thus aiding in the discovery of disease biomarkers and drug targets [699, 700, 701]. Similarly, AI models have revolutionized drug discovery by enhancing various processes, including toxicity prediction, drug release monitoring, Quantitative Structure-Activity Relationship (QSAR) analysis, drug repositioning, physiochemical property prediction, etc [702]. By leveraging RL algorithms, AI-driven drug discovery enhances both virtual screening and de novo drug design. Virtual screening [703, 704, 705] uses RL to identify potential drug candidates from existing compound databases, while de novo design [706, 707, 708, 709] generates novel molecules with optimized biological activity and properties. By integrating RL with active learning and predictive modeling, these systems enhance exploration in vast chemical spaces, reducing experimental costs and improving hit-to-lead timelines [710, 711]. Although they are not fully EmAI systems, they offer valuable insights and core components for the development of future EmAI-driven drug discovery platforms and possess the potential to integrate with physical devices.

### v2

@@ Abstract
keywords:
Embodied artificial intelligence; Multi-modality; Healthcare; Surgical robot; Large language model; Multimodal large language model; World model
@@ 2 Basic AI Techniques for Embodied AI
Figure 3: A detailed breakdown of EmAI core functionalities, with key approaches that support them.
@@ 2.3 High-level Planning
Planning with end-to-end Embodied Large Models. These algorithms stand out by directly mapping high-level instructions to low-level actions, seamlessly integrating perception, planning, and control into a unified system. Recent research shows that these systems leverage deep reinforcement learning and imitation learning to streamline planning and decision-making in complex environments, often surpassing modular pipelines in adaptability and robustness [294, 295]. Notable frameworks such as SayCan [296], PaLM-E [297], and EmbodiedGPT [298] are designed to combine vision encoder embeddings with planning data from LLMs, directly informing the robotic policies for immediate action. However, training embodied end-to-end systems typically requires large-scale datasets. To address this need, simulation data is widely used for its efficiency and safety, though challenges arise from discrepancies in physics, sensors, and real-world complexity. Strategies such as domain randomization [299], domain adaptation [300], and hybrid approaches [301] have been developed to enhance adaptability and bridge the sim-to-real gap, improving real-world performance. In addition, it is also important to design benchmarks to evaluate the embodied decision-making capability among different LLMs [302].
@@ 3.1.1 Pre-Intervention Stage
Recent improvements in EmAI-related technologies for pre-intervention diagnostics and assessments are shaping a new AI-clinician collaboration in the intelligent hospital [356, 357, 358]. EmAI plays various roles in this context, reducing clinicians’ workloads and accelerating diagnostic workflows.
Virtual triage nurse. In modern smart healthcare systems, EmAI-based virtual triage nurses replace human nurses and play a pivotal role in streamlining patient management by directing individuals to the most appropriate clinical departments. These EmAI systems analyze patient-reported symptoms and conduct department ranking based on symptom descriptions [359, 360, 361]. Beyond symptom-based sorting, advanced triage systems integrate wearable health data and EHR to provide a holistic assessment of conditions [362, 363, 364]. In emergency care settings, these systems can even predict patient outcomes and recommend intervention pathways, significantly shortening response times during critical situations [365, 366, 367, 368]. As healthcare systems become more interconnected, intelligent triage systems are increasingly serving as an efficient tool of future infrastructure [364, 369], enabling seamless coordination between primary care, specialist consultations, and hospital admissions.
@@
  • 3.
    Radiology assistant. Interventional Radiology (IR) [419] focuses on using image-guided, minimally invasive techniques to perform diagnostic and therapeutic procedures. The integration of EmAI is further advancing IR by improving procedural accuracy, safety, and operator control in complex medical interventions. For example, robotic systems have applied CT images to improve the accuracy of localization of lesion areas such as pulmonary nodules [107] and navigated placement of thoracolumbar pedicle screws [420] in both preoperative diagnosis and intraoperative procedures. Similarly, robotic systems leveraging X-ray imaging have enhanced cardiac interventions by enabling precise catheter navigation in endovascular catheterization [421, 422]. MRI-guided robotic interventions have facilitated resection margins targeting in procedures like tumor ablation, providing superior visualization of real-time tissue changes during operations to protect critical structures [423, 424, 425].
  • 4.
@@
In summary, fully autonomous ultrasound system [110], human-centric radiology assistant [433], and flexible robotic endoscope system [434] capable of navigating and adjusting diagnostic examinations with minimal human supervision have been developed, achieving a higher level of intelligence and autonomy. Compared to human-led examinations, they can optimize workflows, improve patient care quality, and ensure a safer environment for all involved in the surgical process. Besides, these systems can also function as a module within a more comprehensive EmAI system such as surgical robots, showcasing opportunities to expand their applications.
@@ 3.1.2 In-Intervention Procedure
While the prior discussion centered around the planning of surgical skills into action-level details, within the contemporary medical landscape, there are complex multi-stage surgeries with long-term plans, dividing complex surgeries into multiple sessions for execution. Such surgeries often involve intricate coordination among surgical teams, precise management of patient health over extended periods, and adaptation to unforeseen intraoperative changes. Compared with traditional surgical planning methods, EmAI can enhance multi-stage planning for complex surgeries by integrating various system modules, such as diagnostic tools [395, 396, 457], real-time monitoring systems [458, 459, 460], and predictive analytics [22, 23] for long-term patient health management. These AI-driven modules facilitate the dynamic adaptation of clinical intervention schedules and strategies based on the assessment of patient conditions during successive surgeries [461, 462]. While significant advances have been achieved in individual EmAI modules, seamlessly integrating them into a fully functional surgical planner with global cognition remains a critical focus for future innovation.
@@
Surgical operator. Skillful operators are at the core of every surgical procedure, where precision and timely intervention are crucial. Challenges often arise due to the intricate nature of surgeries and the finite availability of expert operators, which can lead to bottlenecks in patient care. To enhance this, robotic surgery has been a promising trend to offer enhanced precision, dexterity, and minimally invasive approaches to common surgical procedures, leading to improved patient outcomes [463]. In healthcare, a dedicated robot is often more practical and reliable, and has been extensively studied in a variety of settings and shown to provide overall clinical benefits [464].
For example, in laparoscopic surgery, there are robotic systems autonomously performing intestinal anastomoses [466] and rectal cancer resection [467] under human oversight. In retinal microsurgery, comprehensive EmAI systems [468, 469] have been developed to integrate real-time surgical object tracking, segmentation, and model predictive control for intraoperative navigation in delicate microsurgical environments wherein precision is decisive.
@@
For general surgical skills, the potential of robotics has also been widely demonstrated. Robots driven by deep learning and active sensing strategies are trained to use needles through the HOUSTON algorithm [470], which can localize, grasp, and hand over unmodified surgical needles to complete precise manipulation. Their good suturing skill that requires a stable and delicate control system has also been proved by the SmartArm robotic system for a neonatal chest surgery [435]. Similarly, for thread detection and interactive perception, real-time robotic approach [437] is presented for autonomously shortening the tail of a surgical suture using a self-supervised learning framework. In complex surgical scenes, multiple EmAI subsystems may collaborate to integrate multi-source perception and support robotic surgery, requiring a robust control system to direct and coordinate their respective perception-cognition-communication-action loops [471]. Beyond visual perception, other modalities including kinematic data [472], audio data [473], language instructions [102], and tactile perception [474] are integrated to form a holistic view of real-world surgical environments.
@@
Patient digital twin. A patient digital twin represents a detailed, dynamic model of a patient’s biological systems or a part thereof (e.g., anatomy), created using comprehensive and accurate medical data such as imaging studies, physiological measurements, and diagnostic results [114]. It is often combined with advanced visualization devices (e.g., VR/AR) to provide interactive observations, and surgeons can plan, simulate, and optimize surgical pathways based on them [508, 509]. AI-based digital twin creation technology facilitates the identification of surgical patterns [438], prediction of complications or procedural outcomes [509], and generation of medical reports [510, 511]. This technology is often employed to accelerate the learning curve of novice surgeons, helping them familiarize themselves with anatomical structures, surgical procedural contexts, and disease progression [512, 513, 514, 515, 516]. Furthermore, multiple downstream applications have benefited from the surgery digitization process [517]. Clinicians can practice surgical skills and explore human anatomy without relying on real anatomical models using digitalized surgical platforms [518], overcoming the limitation of high costs and scarce training samples. It also offers a digital and accurate platform for robots to train on, improving safety and reliability before deployment in actual surgical procedures [519]. Another approach involves photorealistic surgical images [520] and videos [521] synthesis to benefit the training process of EmAI systems, mitigating challenges related to the high costs and ethical concerns of obtaining surgical data.
@@ 3.2.1 Assistive Robot
The integration of EmAI into healthcare has led to the development of assistive robots that enhance daily care and support of patients. This section focuses on three key applications, including social assistance, daily living assistance, and mobility assistance.
Social guide. In contemporary society, individuals with Autism Spectrum Disorder (ASD), Bipolar Disorder, or other social challenges often face significant barriers in accessing adequate social support. The advancement of EmAI systems, especially those driven by LLMs, has facilitated the development of innovative solutions aimed at addressing these needs [583, 584]. The NAO robot and QTrobot have been utilized to enhance facial recognition and improve eye contact, serving as therapeutic mediators for people with autism and helping them learn to interact with others [585, 586, 587, 588]. Through structured interactions, children can practice social skills such as imitation, turn-taking, engagement, and empathy [589].
@@
Studies have indicated that children with ASD often respond positively to robot-assisted therapy, showing increased engagement and reduced anxiety during sessions [591, 592, 593]. Beyond psychological disorders, EmAI systems like ZORA robots [594, 595, 596] have assisted children with severe physical disabilities in achieving therapeutic and educational goals, as well as helping patients with dementia optimize communication. We present a pipeline for an EmAI system that supports children with social disorders, as shown in Figure 7.
@@ 3.3.1 Rescue Robot
EmAI-powered robots are increasingly developed to serve as human assistants or replacements in various high-risk environments. In rescue scenarios, their potential has been widely explored due to their ability to operate safely in dangerous environments and urgent conditions. In the aftermath of natural disasters or life-threatening emergencies, search and rescue robots aim to provide rapid aid and emergency health services to victims. Thanks to the advancement of recent AI approaches including MLLMs and VLMs, EmAI-equipped robots can efficiently locate and assist survivors while supporting rescue operations [642, 643, 644, 645] in disaster-prone areas, such as earthquake zones. A representative workflow of EmAI-based rescue robots is illustrated in the Figure 8.
@@
The pandemic has also significantly increased the demand for service robots to replace human labor in contaminated areas, particularly for disinfection tasks. This has made efficient and rapid disinfection in hospitals crucial, sparking research into disinfection robots. By integrating mobile robot platforms with advanced disinfection technologies, such as hydrogen peroxide atomization [675] and UVC devices [676, 677], and equipping them with EmAI navigation systems, these robots are able to automate cleaning in complex indoor environments. They are designed to disinfect both floor surfaces and air by eliminating harmful pathogens using a combination of vacuum cleaners, sanitizers, and UV light disinfection methods [678, 679, 680, 681]. By upholding rigorous hygiene standards, these robots significantly mitigate the risk of cross-contamination in healthcare settings.
@@ 3.4 Biomedical Research
EmAI also shows great potential in biomedical and healthcare research [682]. It can synergistically combine human creativity and expertise with AI’s capabilities in handling vast biomedical datasets, navigating complex hypothesis spaces, and executing repetitive experiments. The core functionality of EmAI systems in biomedical research involves planning discovery workflows, performing self-assessments to recognize and address knowledge gaps, and employing structured memory for continual learning [27, 80]. This workflow is illustrated in Figure 9, which shows an automation process of chemical experiments.
@@
Automated lab technician. Experiments are critical parts in biomedical research. EmAI-driven robots automate repetitive tasks, increasing throughput and allowing researchers to focus on more complex analysis tasks. Scientists have created robot chemists capable of autonomously conducting complex chemical reactions and analyses [683, 684]. These robotic systems apply EmAI in laboratory automation, enabling them to design experiments [685, 686, 687], execute procedures [688, 689], and interpret results [690, 691, 692] without human intervention. For instance, a platform leverages AI-based synthesis planning, trained on millions of reactions, to propose synthetic routes, which are then fine-tuned by expert chemists and executed on a robotic flow system [693]. This setup allows for scalable, reproducible synthesis of a variety of compounds, such as pharmaceuticals. Another AI-driven robotic chemist also demonstrated its ability to autonomously perform tasks ranging from synthetic planning to executing experiments in batch reactors [694]. In the field of materials science for photo-catalytic hydrogen production, the robot operates autonomously in a laboratory environment, using Bayesian optimization to navigate a complex experimental space with ten variables, performing 688 experiments over eight days [683]. In summary, EmAI can combine precision handling with autonomous decision-making, allowing it to run experiments continuously without human intervention. These experiment setups minimize human bias, accelerate discovery and production, and enable exploration of large experimental spaces.
@@
Pioneer in drug discovery. The traditional drug discovery process is typically time-consuming and involves substantial costs, presenting significant challenges for pharmaceutical research and development. However, the integration of EmAI into the drug discovery process not only shortens the overall development timeline but also reduces the associated expenses. AI-driven robotic systems can perform rapid screening of vast chemical libraries, identifying potential drug candidates more efficiently than manual methods [695]. In high-throughput screening (HTS) systems, physical components are primarily responsible for automating processes such as sample handling, reaction setup, and data collection through robotics and specialized software for instrument control [696]. By employing liquid handling devices, robotic technology, detectors, and specialized software, researchers have also achieved automation and high efficiency in HTS processes [697] to swiftly evaluate large libraries of compounds for activity against specific biological targets. Additionally, in the analysis and prediction of biological sequences, such as DNA, RNA, or protein structures, LLMs and other approaches are applied to decipher patterns in biological sequences [698]. By treating nucleotides and amino acids as words, they can predict biological structures and functions, thus aiding in the discovery of disease biomarkers and drug targets [699, 700, 701]. Similarly, AI models have revolutionized drug discovery by enhancing various processes, including toxicity prediction, drug release monitoring, Quantitative Structure-Activity Relationship (QSAR) analysis, drug repositioning, physiochemical property prediction, etc [702]. By leveraging RL algorithms, AI-driven drug discovery enhances both virtual screening and de novo drug design. Virtual screening [703, 704, 705] uses RL to identify potential drug candidates from existing compound databases, while de novo design [706, 707, 708, 709] generates novel molecules with optimized biological activity and properties. By integrating RL with active learning and predictive modeling, these systems enhance exploration in vast chemical spaces, reducing experimental costs and improving hit-to-lead timelines [710, 711]. Although they are not fully EmAI systems, they offer valuable insights and core components for the development of future EmAI-driven drug discovery platforms and possess the potential to integrate with physical devices.